Three identical blocks each having a mass m is pushed by a force F on a frictionless table.What is the net force on first block?What force does first block apply on second?What force does second apply on third?Assuming that all the three blocks remain in constant

Grade:11

1 Answers

Saroj Suthar
7 years ago
First draw fbd a=F/3mequation areF-f1 =maf1-f2 =ma f2=maNow add we get the value of a =F/3mFrom aBovey eq we get the value of f1 and f2f1=2F/3 &f2= F/3